# Retrieval-Augmented Semantic Parsing: Using Large Language Models to Improve Generalization

### v1

@@ 6 Conclusion
This paper demonstrates that large language models (LLMs), even without retrieval augmentation, outperform previous encoder-decoder approaches in semantic parsing for Discourse Representation Structures, thereby answering our first research question in the affirmative, setting a new state of the art. We have also shown that our proposed Retrieval-Augmented Semantic Parsing (RASP) framework, which integrates external lexical knowledge, further enhances the performance of LLMs. Notably, RASP nearly doubles the accuracy on out-of-distribution concepts, providing a direct answer to our second research question and confirming its robust generalization ability in open-domain scenarios. Our experiments show that by simply appending relevant lexical information to the model input, the RASP approach offers a practical and intuitive approach that can be easily applied to other meaning representations used in natural language processing, such as AMR Banarescu et al. (2013) and BMR Martínez Lorenzo et al. (2022).
@@ 7 Limitations
The retrieval is proven to provide more information and knowledge to the models. However, retrieval will significantly increase the input length of the model, making it (only) adoptable for the large language models with strong context understanding and long text processing capabilities. Therefore, the RASP framework cannot be directly used to improve previous parsers that rely on other methods, which is also why we only provided results of retrieval-augmented LLMs.
Another limitation is the retrieval source. RASP is based on WordNet, so if a sense is not in WordNet, it will never be guessed. For example, "velvet scooter" (a bird) is not in WordNet, nor is Cobb salad (a dish). Hence, RASP will never make a perfect prediction for such cases. Moreover, the glosses in WordNet, even though carefully crafted by lexicographers in most cases, are sometimes concise, lacking information to separate them from other senses. This makes it difficult for the models to accurately distinguish between different meanings (see Section 5.3).
@@ Appendix A Prompt
Normal Prompt user:

### v2

@@ 6 Conclusion
This paper demonstrates that large language models (LLMs), even without retrieval augmentation, outperform previous encoder-decoder approaches in semantic parsing for Discourse Representation Structures, thereby answering our first research question in the affirmative, setting a new state of the art. We have also shown that our proposed Retrieval-Augmented Semantic Parsing (RASP) framework, which integrates external symbolic knowledge, further enhances the performance of LLMs. Notably, RASP nearly doubles the accuracy on out-of-distribution concepts, providing a direct answer to our second research question and confirming its robust generalization ability in open-domain scenarios. Our experiments show that by simply appending relevant information to the model input, the RASP approach offers a practical and intuitive approach that can be easily applied to other meaning representations used in natural language processing, such as AMR Banarescu et al. (2013) and BMR Mart´ınez Lorenzo et al. (2022).
@@ 7 Limitations
The retrieval process is proven to provide more information and knowledge to the models. However, retrieval will significantly increase the input length of the model, making it (only) adoptable for the large language models with strong context understanding and long text processing capabilities. Therefore, the RASP framework cannot be directly used to improve previous parsers that rely on other methods, which is also why we only provided results of retrieval-augmented LLMs.
Another limitation is the retrieval source. Our implementation of RASP uses WordNet, so if a sense is not in WordNet, it will never be guessed. For example, ”velvet scooter” (a bird) is not in WordNet, nor is Cobb salad (a dish). Hence, RASP will never make a perfect prediction for such cases. Moreover, the glosses in WordNet, even though carefully crafted by lexicographers in most cases, are sometimes concise, lacking information to separate them from other senses. This makes it difficult for the models to accurately distinguish between different meanings (see Section 5.3).
@@ Appendix A Prompt
female.n.02 Name "Johanna" time.n.08 TPR now birdwatch.v.01 Agent -2 Time -1 ELABORATION <1
female.n.02 ANA -3 see.v.01 Experiencer -1 Time +1 Stimulus +3 time.n.08 TPR now
harrier.n.03 kite.n.04 entity.n.01 Sub -2 Sub -1 Sub +1 hobby.n.03
RASP Prompt user: